OpenStudy (anonymous):

If tan x = .67, what is cot x?

hartnn (hartnn):

use tan x = 1/cot x so, cot x = 1/tan x = ...

OpenStudy (anonymous):

1.67?

hartnn (hartnn):

1/0.67 = 1.49

OpenStudy (anonymous):

THANKS!!!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

But how will i find the missing angles if they give me a 90* triangle??

hartnn (hartnn):

whats the entire question ?

hartnn (hartnn):

if one acute angle of a 90 degree triangle is 'x' then other will be 90-x

OpenStudy (anonymous):

Given the following right triangle, what is the measure of the missing side?

hartnn (hartnn):

you need the length of hypotenuse, that is the value of x , right ?